# Multiclass Portfolio Optimization via Variational Quantum Eigensolver with Dicke State Ansatz

### v1

@@ IV Portfolio Optimization and Dicke State
where $a_{i}(\vec{\theta})$ represents amplitude probability as a function of the parameters $\vec{\theta}$. A similar approach was presented by [42], which focused solely on the preparation of a single Dicke state, thus mimicking portfolio optimization without diversification constraints. In this paper, we use the implementation given by [40], creating a non-uniform Dicke state parameterizing the circuit implementation and incorporating the multiclass optimization by including multiple Dicke states representing the multiple classes. Each Dicke state represents a class with a set of products (number of qubits $n$) and will encode the constraints of the exact number of assets which will be selected by class (parameter $k$) (see Fig.3a). The equation below dictates the number of parameters of a Dicke state ansatz and it was derived empirically, for further details see Appendix C
@@ V Results
We addressed the portfolio optimization problem using the SCIP optimizer, treating its solution as the benchmark reference (see AppendixA, Table 4). The running time was determined by averaging the results from $100$ executions. This average running time served as the basis for comparing the performance of classical methods with hybrid VQE routines. However, it is important to note that we did not expect that the VQE approach would outperform the classical methods in terms of speed.
The experiments were performed in three different scenarios as described in Table 2. For Scenario I, we ran the VQE algorithm for $20$ different ansatzes (Dicke state, $R_{y}$, and $18$ variations of Two Local). For Scenarios II and III only the Dicke state was used. For all scenarios, we tested $5$ classical optimizers with $1000$ iterations, $100$ randomly sampled initial points (ansatz parameters), $4096$ shots per circuit, totalizing more than $60$ billion executions ($\approx n_{ansatz}\times n_{optimizers}\times n_{executions}\times n_{shots}\times n_{iterations}$) . More details of the variations of the two-local ansatz used here can be found in Appendix B, Table 5. All data used in this work are publicly available and were obtained using the Yahoo Finance API.
@@ VI Conclusion
Another intriguing aspect we began to evaluate is the relationship between state fidelity and optimal parameters, as discussed in the Appendix E of this paper. In future research, we aim to address the open questions highlighted here and further investigate the use of the Quantum Approximate Optimization Algorithm (QAOA) with Dicke states for multiclass portfolio optimization. This will include exploring different mixer Hamiltonians.
@@
Appendix A Experiment settings
The VQE algorithm simulations were performed using the instance ml.g5.xlarge on Amazon SageMaker. This machine has four virtual CPUs, 16 GB RAM and NVIDIA GPU A10G with 24 GB GDDR6 VRAM. This system runs on Linux, with Python version 3.9.5, CUDA version 12.4, NVIDIA Graphics driver version 550.144.03. The Python libraries used to perform the simulations are: qiskit version 1.3.2 [62], qiskit-aer-gpu version 0.15.1, qiskit-algorithms version 0.3.1, qiskit-optimization version 0.6.1, numpy version 2.0.2, optuna version 4.1.0 and CMA-ES version 0.11.1. All initial points were generated with NumPy random methods using the seed $42$, with the exception for CMA-ES and Random Sampler due to optuna’s parameter sampling method. However, we guarantee that all parameters are in the interval $0,2\pi]$, independently of the method used to sample the parameters values. CMA-ES and Random Sampler were used through the Optuna interface, whereas COBYLA, SPSA and QNSPSA were used through the Qiskit algorithms package.
As mentioned in the beginning of Section V, we solved each scenario proposed with the classical optimizer SCIP and considered its results as a reference. These results are summarized in Table 4.
Scenario Objective function value Running time (s)
I -0.818106 0.01 $\pm$ 0.001
II -1.474237 0.03 $\pm$ 0.001
III -2.00332 0.03 $\pm$ 0.001
Table 4: A summary of the reference values computed using SCIP optimizer for each scenario. The column Objective function value holds the optimal values for each case. The column Running time represents the mean time spent and the standard deviation of 100 runs of the optimization problem.
Appendix B TwoLocal ansatz
The TwoLocal circuit is an ansatz that alternates rotation and entanglement layers. The rotation gates are applied in each qubit individually and the gates $RX$, RYRY, RZRZ, $U_{1}$, $U_{2}$ and $U_{3}$ are some examples of rotation gates that we can use in this ansatz. Meanwhile, the entanglement gates are used to create entanglement between pairs of qubits considered in each entanglement structure design, for instance: we can create entanglement between all qubits or just in some qubits in a way that connects the qubits through a line. In order to create the desired entanglement structure, we can use gates such as: $CX$, CYCY, CZCZ, CRXCRX, CRYCRY and $CRZ$. For instance, in Table 5 we show some TwoLocal settings used in Scenario I experiments.
Name Qubits Parameters Rotation Blocks Entanglement Blocks Entanglement Repetitions Skip Final Rotation Layer
TwoLocal1 10 30 3 x RY CX Full 1 True
TwoLocal2 10 30 1 x RY CX Full 2 True
TwoLocal3 10 30 1 x RY CX Full 3 False
TwoLocal4 10 30 3 x RY CX Linear 1 True
TwoLocal5 10 30 1 x RY CX Linear 2 True
TwoLocal6 10 30 1 x RY CX Linear 3 False
TwoLocal7 10 30 3 x RY CX Reverse Linear 1 True
TwoLocal8 10 30 1 x RY CX Reverse Linear 2 True
TwoLocal9 10 30 1 x RY CX Reverse Linear 3 False
TwoLocal10 10 30 3 x RY CX Circular 1 True
TwoLocal11 10 30 1 x RY CX Circular 2 True
TwoLocal12 10 30 1 x RY CX Circular 3 False
TwoLocal13 10 30 3 x RY CX SCA 1 True
TwoLocal14 10 30 1 x RY CX SCA 2 True
TwoLocal15 10 30 1 x RY CX SCA 3 False
TwoLocal16 10 30 3 x RY CX Pairwise 1 True
TwoLocal17 10 30 1 x RY CX Pairwise 2 True
TwoLocal18 10 30 1 x RY CX Pairwise 3 False
Table 5: Settings used to compose each TwoLocal variant consired in the experiments in Scenario I. All circuits for Scenario I were created to have 30 parameters and 10 qubits, then we just varied the amount of rotation blocks, entanglement structure, the number of layers repetitions and whether we skip the final rotation layer. In this case we only used RYRY and CXCX gates because relative phases are not relevant.
Appendix C Number of parameters equation
In this appendix, we will give details about how we derived the equation 6 empirically. The total number of parameters is directly related to the number of $CRY$ and $CCRY$ gates, and each of these gates has one free parameter. We defined a range of values for $n$ and considered [1,n1][1,n-1] as the range for $k$. Then, we constructed the Dicke state ansatz for all values of $n$ and $k$ defined in the ranges mentioned above, and computed the total number of parameterized gates. With these data in our hands, we were able to plot the curves (see Figure 8) and also perform a regression to derive a formula that dictates the relationship between the number of qubits and the Hamming weight, with the number of parameters in Dicke state ansatz.
Considering a fixed value for $k$ and varying the number of qubits $n$ in the range $[k+1,d]$, where $d>k+1$, and building the Dicke state ansatz circuit of each value of $n$, we can calculate the number of parameters. Plotting $n_{p}$ by $n$, we saw a polynomial function of degree 1, then we can write the following expression
$n_{p}=mn+b.$ (8)
The angular coefficient can be calculated through
$m=\frac{n_{p}^{f}-n_{p}^{i}}{n_{f}-n_{i}},$ (9)
and computing the value of $m$ with the data obtained from the quantum circuits of Dicke state ansatz, we found that $m=k$, what implies that the angular coefficient is equal to the Hamming weight of the Dicke state. The linear coefficient can be computed using the following expression
$b=n_{p}^{i}-kn_{i},$ (10)
and in order to find out if there is a general rule for $b$, we tested different values of $k$ for a fixed $d$ and we found that the value of the linear coefficient is given by
$b=\frac{k(k+1)}{2}.$ (11)
Thus, combining all the elements computed above we obtain the following equation for a Dicke state ansatz with $n$ qubits and Hamming weight $k$
$n_{p}=kn-\frac{k(k+1)}{2}.$ (12)
Figure 8: Scaling of the number of parameters in Dicke state ansatz for different numbers of qubits and varying the Hamming weight $k$. Note that for $k=0$ and $k=n$ we don’t need a parameterized Dicke state, because $k=0$ is equivalent to $|0\rangle^{\otimes n}$ and $k=n$ is the state $|1\rangle^{\otimes n}$, then the range applied to $k$ is [1,n1][1,n-1]. The curves defined by different values of $n$ and $k$ can be unified by the expression nkk(k+1)/2nk-k(k+1)/2, as depicted in the plot in the right.
Appendix D Dicke state ansatz on noisy devices
Once a good ansatz has been identified, it is important to improve its implementation in hardware to improve the results. In reference [42], the authors proposed a Dicke state ansatz version that is more efficient to run in the current noisy quantum hardware, reducing the number of entangling gates and also considering qubits connections constraints. Another way to improve the depth of Dicke state circuit proposed in [42, 40], it is the parallelization of entangling gates, with that we can optimize circuit depth aiming to create quantum circuits which are suitable for current quantum hardware coherence time. A topic to highlight is the Dicke state sensitivity to errors of two types: bit-flip and readout errors. For instance, we can measure the impact of bit-flips, in our ansatz, modeling this type of error as described below and manipulating the probability $p$ of occurrence of a bit-flip [63]
$\rho=(1-p)\rho_{0}+p\sigma_{x}\rho_{0}\sigma_{x},$ (13)
where $\rho_{0}$ represents the density matrix related with the Dicke state ansatz. Then we can use the noisy state $\rho$ to calculate the Hamiltonian expected value through [64]
$\langle H\rangle=\textrm{tr}(\rho H),$ (14)
and with that we will be able to estimate the impact on our objective function. Both errors can destroy the main characteristics of this quantum state, that is, a constant Hamming weight, because depending of the number of bit-flips and in which qubits they will occur, this kind of error can increase or reduce the Hamming weight, which will reflect in adding Hamiltonian terms that will have a great impact on Hamiltonian expectation value. Then, it is expected that the results obtained from current noisy quantum hardware will not achieve the same level of convergence that we achieved with noiseless simulation, because bit-flips and readout errors can make parameter optimization more challenging, since noise can also induce Barren Plateaus [65].
One last type of error that can cause issues during parameter optimization is the coherent error, more specifically, an error of the form $U(\theta+\delta\theta)$, where $\delta\theta$ represents a displacement in the original parameter. In order to mitigate the damage that the errors described above can cause, we can use some error mitigation techniques. For example, to mitigate errors in expectation values we can use Zero Noise Extrapolation (ZNE) [66, 67, 68], Probabilistic Error Cancellation (PEC) [69, 70] and Probabilistic Error Amplification (PEA) [71, 72]. For readout errors, we can apply statistical corrections [73, 74], M3 technique [73] or Twirled Readout Error Extinction (T-REX) [74]. The usage of the error mitigation techniques mentioned above, combined with advanced transpiling methods are highly recommended, in order to produce optimized quantum circuits to current hardware, improving the chances of obtaining good results. Another interesting direction that we would like to explore is the usage of Conditional Value at Risk (CVaR) as the objective function [75, 76], as CVaR seems to be a noise-resistant loss function for variational quantum algorithms.
Appendix E Quantum state fidelity analysis
Figure 9: A visual analysis of the state fidelity behavior in each scenario proposed. The heatmap represents the state fidelity between of each experiment with Dicke state and CMA-ES, then each point reflects the similarity between the quantum states produced by two sets of parameters obtained at the end of the corresponding experiments. The main diagonal is always equal to 1, because it is the fidelity between the quantum states defined by the same set of parameters. The colored histogram in the Y-axis shows how the points in the heatmap are distributed in the state fidelity scale. The gray histogram represents the state fidelity distribution obtained when we compute the fidelity between the quantum states produced by the optimal set of parameters for each experiments and the reference state for each case: (a) Scenario I results. (b) Scenario II results. (c) Scenario III results.
An additional analysis that we performed was on the quantum state fidelity between states at the end of the optimization process, but also on their state fidelity with reference state. Here we only consider results obtained with Dicke state ansatz simulations using CMA-ES as optimizer, these results are summarized in Figure 15.
State fidelity computation was done considering pure states, then we used the following mathematical expression [63]
$F(|\psi(\vec{\theta^{*}_{i}})\rangle,|\psi(\vec{\theta^{*}_{j}})\rangle)=|\langle\psi(\vec{\theta^{*}_{j}})|\psi(\vec{\theta^{*}_{i}})\rangle|^{2},$ (15)
where $\vec{\theta^{*}_{i}}$ and $\vec{\theta^{*}_{j}}$ represent two different optimal set of parameters for the states described in Table 2. To compute the state fidelity, we consider a parameterized quantum circuit that is used to compute the kernel in QSVM [77], where we apply an unitary operator $U(\vec{\theta})$ that prepares the desired state and its adjoint with different set of parameters, as stated in the following expression
$|\Phi\rangle=U^{\dagger}(\vec{\theta_{j}})U(\vec{\theta_{i}})|0\rangle^{\otimes\ n}.$ (16)
Then, our estimation of fidelity will be equal to the probability of measuring $|0\rangle^{\otimes\ n}$. This fidelity estimation could also be done with the SWAP test [78].
Our finding here was that these different sets of optimal parameters lead to quantum states that have a high fidelity to the reference state. We also noted a certain level of similarity between these quantum states as shown by the heatmaps in Figure 15. However, we cannot extract useful information from the distance between the parameter sets, because it is not clear how we can define what is close or distant in the parameter space.
Perhaps, with this information in hand, we can extract some useful information about the optimization landscape to improve the success of this method. A future research line would design a neural network to provide a good initial starting set of parameters considering this information about the similarity between the states that have a high fidelity with the reference state and are similar between them.

### v2

@@ IV Portfolio Optimization and Dicke State
where $a_{i}(\vec{\theta})$ represents amplitude probability as a function of the parameters $\vec{\theta}$. A similar approach was presented by [42], which focused solely on the preparation of a single Dicke state, thus mimicking portfolio optimization without diversification constraints. In this paper, we use the implementation given by [40], creating a non-uniform Dicke state parameterizing the circuit implementation and incorporating the multiclass optimization by including multiple Dicke states representing the multiple classes. Each Dicke state represents a class with a set of products (number of qubits $n$) and will encode the constraints of the exact number of assets which will be selected by class (parameter $k$) (see Fig.3a). The equation below dictates the number of parameters of a Dicke state ansatz and it was derived empirically, for further details see Appendix XII
@@ V Results
We addressed the portfolio optimization problem using the SCIP optimizer, treating its solution as the benchmark reference (see AppendixX, Table 4). The running time was determined by averaging the results from $100$ executions. This average running time served as the basis for comparing the performance of classical methods with hybrid VQE routines. However, it is important to note that we did not expect that the VQE approach would outperform the classical methods in terms of speed.
The experiments were performed in three different scenarios as described in Table 2. For Scenario I, we ran the VQE algorithm for $20$ different ansatzes (Dicke state, $R_{y}$, and $18$ variations of Two Local). For Scenarios II and III only the Dicke state was used. For all scenarios, we tested $5$ classical optimizers with $1000$ iterations, $100$ randomly sampled initial points (ansatz parameters), $4096$ shots per circuit, totalizing more than $60$ billion executions ($\approx n_{ansatz}\times n_{optimizers}\times n_{executions}\times n_{shots}\times n_{iterations}$) . More details of the variations of the two-local ansatz used here can be found in Appendix XI, Table 5. All data used in this work are publicly available and were obtained using the Yahoo Finance API.
@@ VI Conclusion
Another intriguing aspect we began to evaluate is the relationship between state fidelity and optimal parameters, as discussed in the Appendix XIV of this paper. In future research, we aim to address the open questions highlighted here and further investigate the use of the Quantum Approximate Optimization Algorithm (QAOA) with Dicke states for multiclass portfolio optimization. This will include exploring different mixer Hamiltonians.
@@
Acknowledgements. We acknowledge the Instituto de Ciência e Tecnologia Itaú for all the support during this research project.
@@
VII Data Availability
The datasets generated and/or analysed during the current study are available in the Yahoo Finance and can be downloaded using the following python library: yfinance. The period of the datasets considered in this work are: 2023-12-18 until 2024-12-16 and 2024-01-02 until 2025-01-02.
VIII Code Availability
The underlying code for this study is not publicly available but may be made available to qualified researchers on reasonable request from the corresponding author.